Consider the linear system X' = AX where A is defined by

A = \begin{bmatrix}a &-b \\b & a \end{bmatrix},

where a and b are real numbers. Assume that the determinant of A is not zero.

Classify the equilibrium solution (0,0) depending on the signs of a and b. Also, sketch a few trajectories for each case, including a few tangent vectors.
Answer #1

X AX با ۔ A ده Now, find characterstic Eqni d²- (ataildt (at i=0 (.. deta= a + b to). andt a+6= Il azt. 6² t either cito oNote in this Seal. case both eigen, value are Now. thin a² ca+ 6 yo ux-ve may in this be tve crise eigen: valice complex Conthus -62do..not possible. so Characterstic Ea? never have a non-zeto distinct seal sout 8o. Equilibtium fol. is node for any